OpenStudy (anonymous):

What is the probability that a randomly selected 2 digit number is prime?

OpenStudy (anonymous):

Please explain that please....

OpenStudy (anonymous):

Wait, nevermind, that's co-primality, the density of primes is \[ \pi(x)\sim \frac{\ln x}{x} \]

OpenStudy (anonymous):

I don't understand this

hartnn (hartnn):

First u need to find how many two digit prime numbers are there.(say x) Total 2 digit numbers=90,right? so your probability=x/90 make sense?

OpenStudy (anonymous):

2/90?

OpenStudy (anonymous):

It's not right, that answer isn't there in my multiple choice sheet

hartnn (hartnn):

no! there are many prime numbers of 2 digits,not only 2...

OpenStudy (anonymous):

Yeah, this proof is absolutely ridiculous, but, if we write a quick computer program to solve it for us (the only human way I see of being able to do this nicely), we get 23.33%

OpenStudy (anonymous):

Which is the total number of primes that are two digits over the total amount of two-digit numbers. As @hartnn stated.

OpenStudy (anonymous):

Since there are 21 two-digit primes.

OpenStudy (anonymous):

Oops there are about 20 prime numbers

OpenStudy (anonymous):

OKAY I GOT THE ANSWER NOW

OpenStudy (anonymous):

:D 7/30 right?

hartnn (hartnn):

yup :)

OpenStudy (anonymous):

But aren't there from only 99-10 two digit numbers? So how is it over 90?

OpenStudy (anonymous):

There are 99-10+1 two-digit numbers (remember 10 is also two digits!)